# Towards Heterogeneous Continual Graph Learning via Meta-knowledge Distillation

### v1

@@ 5 Experiments
Figure 2: Performance comparison on the Yelp dataset under two settings using HAN as the backbone model.
Ablation Study To evaluate the effectiveness of individual modules in MKD, we conducted comprehensive ablation experiments using HAN as the backbone network across four benchmark datasets. The results are presented in Table 2.
We first examined the impact of each component by sequentially removing Experience Replay (ER), Heterogeneity-aware Knowledge Distillation (HKD), and the Gradient-based Meta-learning Module (GMM) while keeping other components intact. The results demonstrate that removing any module leads to performance degradation in both AP and AF, indicating each component’s essential role in our framework. Notably, the removal of GMM shows relatively smaller performance impact (0.3-3.4% decrease in AP compared to 1.5-11.0% for ER/HKD), which aligns with its primary function of facilitating rapid adaptation to new tasks rather than long-term knowledge retention - the latter being mainly handled by the synergistic effect of ER and HKD. Particularly, the absence of HKD causes the most significant performance deterioration (3.6-12.0% increase in FR), highlighting its crucial role as the core mechanism against catastrophic forgetting.
Furthermore, we replaced our proposed Efficient Heterogeneous Subgraph Sampling (E-HSS) with the basic Coverage Maximization (CM) strategy. As shown in the last row of Table 2, E-HSS demonstrates substantial performance advantages, especially on edge-dense datasets like Yelp. These results confirm that E-HSS significantly outperforms conventional sampling methods in preserving the original topological structure of heterogeneous graphs. We further investigate the impact of each component in the HKD module in Appendix C.1.
@@
Table 2: Ablation results of MKD using HAN as the backbone. We remove Experience Replay (ER), Heterogeneity-aware Knowledge Distillation (HKD), and the Gradient-based Meta-learning Module (GMM) individually. And “w/o E-HSS” denotes replacing the proposed Efficient Heterogeneous Subgraph Sampling with Coverage Maximization applied to all node types.
@@
w/o E-HSS 92.4 $\pm$ 1.4 6.0 $\pm$ 3.1 75.4 $\pm$ 2.2 12.1 $\pm$ 5.0 56.2 $\pm$ 0.4 18.0 $\pm$ 0.5 45.7 $\pm$ 14.0 23.7 $\pm$ 15.6
Hyperparameters Analysis We analyze the impact of key hyperparameters on model performance, including the loss weight ($\lambda_{kd}$) in knowledge distillation (KD), the sampling budget for experience replay, and the shot number (i.e., the number of sampled nodes used in meta-learning), as shown in Figure 3. We further discuss in Appendix C.2 the impact of the trade-off weight used to balance the losses of two submodules in the knowledge distillation component. The experimental findings are as follows:
@@
Distillation loss Weight ($\lambda_{kd}$): On IMDB, higher weights better support knowledge transfer, while lower values ($\lambda_{kd}<0.2$) lead to increased forgetting. On Freebase, smaller weights ($0.1\sim 0.3$) help mitigate forgetting, but larger weights ($\lambda_{kd}>0.3$) hinder adaptability. For Yelp, the 3-way setting introduces large task shifts, requiring higher weights for stability and effective forgetting mitigation. Sample Budget: The number of sampled nodes influences the stability-plasticity trade-off. Too few nodes lead to insufficient information, degrading performance, while excessive sampling limits the model’s adaptability to new tasks. As shown in Figure 3, on datasets like Yelp, a large sampling budget (e.g., >20) significantly impairs learning on new tasks due to the accumulation of historical information. Shot Number: The number of sampled nodes impacts the stability-plasticity trade-off. Meta-learning enables rapid adaptation to new tasks. As shown in the results, model performance steadily improves up to 10-shot, but declines at 20-shot, likely due to overfitting to the current task’s data patterns. The optimal shot number is 10, balancing historical information retention and computational efficiency.
@@
Figure 3: Sensitivity analysis of MKD with respect to key hyperparameters. The shaded areas represent variances. From left to right: (1) distillation loss weight, (2) experience replay sampling budget, and (3) shot number used in meta-learning. Due to space constraints, we only report the Average Performance (AP) here.
6 Conclusion
Our proposed method integrates meta-learning, heterogeneity-aware knowledge distillation, and experience replay, specifically designed to address challenges in heterogeneous continual graph learning. By jointly considering the diversity within target node types and the structural relations across different node types, our efficient sampling strategy preserves critical topological and semantic information with limited data. The gradient-based meta-learning module enables rapid adaptation to new tasks, while the knowledge distillation module aligns knowledge across tasks at both prediction and semantic levels, effectively balancing model stability and plasticity. Extensive experiments on multiple real-world datasets validate the effectiveness of the proposed method on the Heterogeneous Continual Graph Learning (HCGL) problem.
7 Acknowledgment
@@
Appendix A Supplement Methodology
A.1 Detailed Algorithm
The detailed framework of the proposed method MKD for HCGL is described in Algorithm 1.
Algorithm 1 Framework of our MKD method
0:  Continual tasks $\mathcal{T}=\{\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{T}\}$; Buffer set $\{\mathcal{B}_{\tau}\}_{\tau\in\mathcal{A}}$; Learning rate $\alpha$ for the inner update; Learning rate $\beta$; Number of target type nodes in each class added to $\mathcal{B}$: e.
0:  Model $f_{\theta}$ which can mitigate catastrophic forgetting of preceding tasks.
1:  Initialize model parameters $\theta$
2:  while Continual Tasks $\mathcal{T}$ remains do
3:     Obtain current training set $V^{tr}_{t}$
4:     Obtain teacher models $\{f_{T}\}$ and Buffer set $\{\mathcal{B}_{\tau}\}_{\tau\in\mathcal{A}}$
5:     Select meta-learning examples $\mathcal{E}=Select(\mathcal{V}_{i}^{tr},e)$
6:     for $epoch=1$ to $E$ do
7:        Update the parameters $\theta$ using $\mathcal{E}$ via Equation 1
8:        Replay the experience via Equation 7
9:        Distill the soft knowledge and semantic information from teacher models via Equation 9 and 10
10:        Update the parameters of student model by optimizing Equation 12
11:     end for
12:     Select target type nodes via 2 and other types nodes via 5
13:     Add all sampled nodes to Buffer
14:     Add current task model to teacher models $\{f_{T}\}$
15:     $\mathcal{T}=\mathcal{T}\setminus\{\mathcal{T}_{i}\}$
16:  end while
17:  return  Model $f_{\theta}$
A.2 Extension to General HGNNs
@@
$e_{ij}=\left(h_{i}^{(L-1)}W^{(L)}\right)^{\top}\tanh\left(h_{j}^{(L-1)}W^{(L)}\right),$ (14)
@@
$\boldsymbol{\alpha}_{i}=\text{softmax}(e_{i1},e_{i2},\dots,e_{i|V_{\tau_{t}}|})$ (15)
@@
$\mathcal{L}_{\text{sem}}=\sum_{i\in V_{\tau_{t}}}\left\|\boldsymbol{\alpha}_{i}^{(T)}-\boldsymbol{\alpha}_{i}^{(S)}\right\|_{2}$ (16)
@@
$\mathcal{L}_{\text{sem}}=\sum_{r\in\mathcal{R}}\left\|\boldsymbol{\beta}_{r}^{(T)}-\boldsymbol{\beta}_{r}^{(S)}\right\|_{2}$ (17)
@@ Appendix B Supplemental experiment setups
B.1 Details of dataset
DBLP is a bibliographic dataset in the field of computer science. A widely adopted subset is used, containing four research areas, where nodes represent authors, papers, terms, and venues. IMDB is a dataset derived from a movie information platform. A subset including classes such as Action, Comedy, Drama, Romance, and Thriller is selected for use. Freebase is a large-scale knowledge graph. A subgraph comprising eight genres of entities and approximately one million edges is sampled following procedures outlined in prior studies. Yelp contains a network constructed from businesses, users, locations, and reviews. Although node features are not available, many business nodes are annotated with one or more labels from sixteen predefined categories.
Table 3: Details of datasets and continual learning tasks setting
@@ B.1 Details of dataset
B.2 Baseline Introduction
@@
Elastic weight consolidation (EWC) [21] is a technique that regularizes the loss function, such that the model is encouraged to only modify the weights that are less important for the previous tasks. This is achieved by penalizing changes to weights that have large importance for the previously learned tasks, thereby mitigating catastrophic forgetting when learning new tasks.
Memory aware synapses (MAS) [20] is a regularization-based method that measures the importance of parameters based on the sensitivity of predictions to these parameters. When learning a new task, the model adjusts the weights according to the network’s activations to update the important parameters relevant to the new task data.
Topology-aware weight preserving(TWP) [25] is a method introduced for graph continual learning that incorporates weight preservation mechanisms based on graph topology. By considering the local structure of the graph, TWP aims to overcome catastrophic forgetting. After computing the loss, the importance scores of the model weights are calculated according to the topology of the given graph, and the loss is regularized accordingly. This method leverages the properties of the graph.
ER-GNN [26] stores samples from previous tasks as experiences and replays them when learning new tasks. After learning task $T_{i}$, sample nodes are saved to the buffer using a selection function. When learning the next task $T_{i+1}$, separate graphs are constructed for each learned task $\{T_{t}\}_{t=1,\dots,i}$. Then, the GNN is trained using these graphs. The overall loss is calculated by regularizing the current task’s loss with the loss from the separately constructed graphs through experience replay.
MetaCLGraph [28] combines experience replay and meta-learning. In this method, the initial parameters of the model are calculated using the current task data. When learning a new task, stored samples from previous tasks are merged with the current task data to form a new graph, which is then used to update the model parameters.
FTF-ER [57] combines feature and global topological information by normalizing and weightedly integrating two types of node importance scores to evaluate node importance. Specifically, it introduces the Hodge Potential Score (HPS) module to capture global topological information. When learning a new task, a subgraph is induced using all the experience nodes stored in the buffer, the overall loss is calculated as the sum of the loss for the current task and the loss of the subgraph.
B.3 Experimental Setting
@@
Figure 4: Illustration of the domain incremental setting.
B.4 Implementation Detail
To evaluate the effectiveness and generalizability of our method, we conduct experiments on three HGNNs backbones: HAN [6], MAGNN [4], and HGT [13]. Adam optimizer is used and the initial learning rate is set to 0.005 for all datasets. For HAN, each task is trained for 200 epochs with an early stopping patience of 100. For MAGNN, each task is trained for 100 epochs, with the early stopping patience set to 5 on DBLP and 10 on IMDB. For HGT, all tasks are trained for 300 epochs with a patience of 30. For datasets trained with mini-batches, the batch size is set to 8. We adopt coverage maximization as the selection function for all experience-replay-based continual learning methods. The buffer size is uniformly set to 50 for the target node type in experience replay baselines (e.g., ER-GNN, MetaCLGraph, and FTF-ER), and 20 under mini-batch settings. For non-target node types, the buffer size is set to 200. The hyperparameters for regularization-based baselines (e.g., EWC, MAS, and TWP) are set following the TWP official repository and the benchmark study [24].
@@ Appendix C Additional Results and Analysis
C.1 Additional Ablation Study

### v2

@@ 5. Experiments
Figure 3. Performance comparison on the Yelp dataset under two settings using HAN as the backbone model.
Performance Evaluation
The overall performance across three heterogeneous graph benchmarks is summarized in Table 1. Additional insights from task-wise accuracy retention on Yelp are shown in Figure 3 and Figure 4. We highlight the following key observations: (1) All HGNNs exhibit catastrophic forgetting on previous tasks. For example, on the DBLP dataset, the average forgetting (AF) for HAN and MAGNN is over 26% and 34%, respectively; on the IMDB dataset, the AF for all HGNNs is over 26%, 17%, and 21%, respectively. (2) Across all three HGNN backbones, HERO substantially outperforms state-of-the-art continual graph learning methods in terms of average performance (AP) while maintaining the lowest forgetting (AF). For example, on DBLP with HAN backbone, HERO improves AP by +2.7% over the strongest baseline (MAS) while reducing AF to 4.3%, close to the upper bound of joint training. Similar consistent performance gains are observed on IMDB and Freebase. These results confirm the effectiveness of HERO as a holistic framework for balancing knowledge adaptation and retention in HCGL. Furthermore, HERO demonstrates relatively low variance across the 5 runs, indicating its ability to consistently perform well in various scenarios. (3) It is noteworthy that, in some cases, memory-replay based methods (e.g., FTF-ER, MetaCLGraph, and ER-GNN) perform better than HERO in terms of AF, despite their lower AP scores compared to HERO. This is because these methods rely on retraining nodes from previously learned tasks. Once enough nodes are sampled, these models can largely mitigate catastrophic forgetting. However, this also sacrifices performance when learning new tasks. Although these methods attempt to address the new-old task trade-off through different strategies (e.g., ER-GNN reweights the loss between new and old tasks), they still cannot fully avoid the performance drop. We further discuss this in Appendix C.4. (4) We also observe a notable phenomenon: in the 3-way setting, memory-replay methods generally beat regularization-based methods, while in the 2-way setting the opposite holds. Further analysis of Figure 4 shows that regularization methods are less effective than memory-replay at overcoming forgetting of earlier tasks, but they achieve better performance on recent tasks, which is the “performance sacrifice” effect mentioned earlier. Memory-replay tends to accumulate redundant or conflicting information over long task sequences, which can harm adaptation to recent tasks. HERO mitigates this by filtering redundant information and applying knowledge distillation, preserving historical knowledge while maintaining stronger performance on recent tasks.
Figure 4. Visualization: Accuracy matrices on the Yelp dataset under 2-way setting.
@@
Table 2. Ablation results of HERO using HAN as the backbone. We remove Experience Replay (ER), Heterogeneity-aware Knowledge Distillation (HKD), and the Gradient-based Meta-learning Module (G-MM) individually. And “w/o DiSCo” denotes replacing the proposed Diversity Sampling with semantic Consistency with Coverage Maximization applied to all node types.
@@
w/o DiSCo 92.4 $\pm$ 1.4 6.0 $\pm$ 3.1 75.4 $\pm$ 2.2 12.1 $\pm$ 5.0 56.2 $\pm$ 0.4 18.0 $\pm$ 0.5 45.7 $\pm$ 14.0 23.7 $\pm$ 15.6
Ablation Study
To evaluate the effectiveness of individual modules in HERO, we conducted comprehensive ablation experiments using HAN as the backbone network across four benchmark datasets. The results are presented in Table 2.
We first examined the impact of each component by sequentially removing Experience Replay (ER), Heterogeneity-aware Knowledge Distillation (HKD), and the Gradient-based Meta-learning Module (G-MM) while keeping other components intact. The results demonstrate that removing any module leads to performance degradation in both AP and AF, indicating each component’s essential role in our framework. Notably, the removal of G-MM shows relatively smaller performance impact (0.3-3.4% decrease in AP compared to 1.5-11.0% for ER/HKD), which aligns with its primary function of facilitating rapid adaptation to new tasks rather than long-term knowledge retention, the latter being mainly handled by the synergistic effect of ER and HKD. Particularly, the absence of HKD causes the most significant performance deterioration (3.6-12.0% increase in FR), highlighting its crucial role as the core mechanism against catastrophic forgetting.
Furthermore, we replaced our proposed Diversity Sampling with semantic Consistency (DiSCo) with the basic Coverage Maximization (CM) strategy. As shown in the last row of Table 2, DiSCo demonstrates substantial performance advantages, especially on edge-dense datasets like Yelp. These results confirm that DiSCo significantly outperforms conventional sampling methods in preserving the original topological structure of heterogeneous graphs. We further investigate the impact of each component in the HKD module in Appendix C.2.
Hyperparameters Analysis
@@
We analyze the impact of key hyperparameters on model performance, including the loss weight ($\lambda_{kd}$) in knowledge distillation (KD), the sampling budget for experience replay, and the shot number (i.e., the number of sampled nodes used in meta-learning), as shown in Figure 5. We further discuss in Appendix C.3 the impact of the trade-off weight used to balance the losses of two submodules in the knowledge distillation component. The experimental findings are as follows: Distillation loss Weight ($\lambda_{kd}$): On IMDB, higher weights better support knowledge transfer, while lower values ($\lambda_{kd}<0.2$) lead to increased forgetting. On Freebase, smaller weights ($0.1\sim 0.3$) help mitigate forgetting, but larger weights ($\lambda_{kd}>0.3$) hinder adaptability. For Yelp, the 3-way setting introduces large task shifts, requiring higher weights for stability and effective forgetting mitigation. Sample Budget: The sample budget influences the stability-plasticity trade-off. Too few nodes lead to insufficient information, degrading performance, while excessive sampling limits the model’s adaptability to new tasks. As shown in Figure 5, on datasets like Yelp, a large sampling budget (e.g., ¿20) significantly impairs learning on new tasks due to the accumulation of historical information. Shot Number: The number of sampled nodes impacts the stability-plasticity trade-off. Meta-learning enables rapid adaptation to new tasks. As shown in the results, model performance steadily improves up to 10-shot, but declines at 20-shot, likely due to overfitting to the current task’s data patterns. The optimal shot number is 10, balancing historical information retention and computational efficiency.
@@
Figure 5. Sensitivity analysis of HERO with respect to key hyperparameters. The shaded areas represent variances. From left to right: (1) distillation loss weight, (2) experience replay sampling budget, and (3) shot number used in meta-learning. Due to space constraints, we only report the Average Performance (AP) here.
Scalability Analysis
Based on the HAN backbone, we evaluate recent continual graph learning approaches and proposed HERO framework on four web graph datasets using a single RTX 3080Ti GPU. Table 3 reports the average training time per task. The results show that the regularization-based method TWP is clearly more efficient than replay-based methods in training, highlighting an open problem that we aim to address in future work. Nevertheless, our method significantly outperforms other replay-based approaches, demonstrating substantial improvements in computational efficiency over traditional replay-based methods. Although TWP is slightly more efficient than HERO, it consistently underperforms in terms of accuracy and forgetting across all benchmarks. Considering HERO’s strong empirical performance, we argue that HERO provides a practical balance between efficiency and effectiveness, which can also be effectively trained on large-scale and long-sequence HCGL tasks. Due to length limitation, we provide results of inference time in Appendix C.1.
Table 3. Training time (s) comparison on four datasets. Lower is better. Numbers in parentheses denote the ratio relative to HERO.
Method DBLP IMDB Freebase Yelp
TWP 0.1566 (x0.42) 0.0320 (x0.78) 0.0676 (x0.59) 3.0195 (x0.50)
FTF-ER 1.6193 (x4.39) 0.2382 (x5.78) 0.3514 (x3.06) 8.4238 (x1.39)
ER-GNN 0.4105 (x1.11) 0.0533 (x1.29) 0.0780 (x0.68) 6.6517 (x1.10)
MetaCLGraph 1.7702 (x4.80) 0.3101 (x7.53) 0.6144 (x5.34) 9.7515 (x1.61)
HERO (Ours) 0.3689 (x1.00) 0.0412 (x1.00) 0.1150 (x1.00) 6.0433 (x1.00)
6. Conclusion
In this work, we systematically study the problem of heterogeneous continual graph learning (HCGL), a critical yet underexplored challenge in dynamic web environments such as recommendation, knowledge graphs, and social networks. We present HERO, a heterogeneity-aware continual learning graph framework designed for evolving web graphs. HERO introduces two key innovations: (1) DiSCo (Diversity Sampling with semantic Consistency), a sampling strategy that maximizes node diversity and expands subgraphs along metapaths to preserve important semantic and structural information. This design enables HERO to achieve promising performance with a significantly smaller buffer size compared to existing replay-based approaches, highlighting its memory efficiency. (2) Heterogeneity-aware Distillation, which aligns semantics and representation across tasks, ensuring consistent knowledge transfer. By integrating these components within a holistic mechanism, HERO effectively balances stability and plasticity, achieving superior performance and scalability across multiple web graph benchmarks.
7. Acknowledgment
@@
Appendix A Extension to General HGNNs
@@
(14) $e_{ij}=\left(h_{i}^{(L-1)}W^{(L)}\right)^{\top}\tanh\left(h_{j}^{(L-1)}W^{(L)}\right),$
@@
(15) $\boldsymbol{\alpha}_{i}=\text{softmax}(e_{i1},e_{i2},\dots,e_{i|V_{\tau_{t}}|})$
@@
(16) $\mathcal{L}_{\text{sem}}=\sum_{i\in V_{\tau_{t}}}\left\|\boldsymbol{\alpha}_{i}^{(T)}-\boldsymbol{\alpha}_{i}^{(S)}\right\|_{2}$
@@
(17) $\mathcal{L}_{\text{sem}}=\sum_{r\in\mathcal{R}}\left\|\boldsymbol{\beta}_{r}^{(T)}-\boldsymbol{\beta}_{r}^{(S)}\right\|_{2}$
@@ Appendix B Supplemental experiment setups
B.1. Details of dataset
DBLP is a bibliographic dataset in the field of computer science. A widely adopted subset is used, containing four research areas, where nodes represent authors, papers, terms, and venues.
IMDB is a dataset derived from a online movie information platform. A subset including classes such as Action, Comedy, Drama, Romance, and Thriller is selected for use.
Freebase is a large-scale knowledge graph originally constructed for web search and semantic applications. A subgraph comprising eight genres of entities and approximately one million edges is sampled following procedures outlined in prior studies.
Yelp is a review network constructed from businesses, users, locations, and reviews. Although node features are not available, many business nodes are annotated with one or more labels from sixteen predefined categories. As a prototypical web application graph, Yelp embodies real-world challenges in recommendation and user–business interaction modeling.
Table 4. Details of datasets and continual learning tasks setting
@@ B.1. Details of dataset
Figure 6. Illustration of the domain incremental setting.
B.2. Baseline Introduction
@@
Elastic weight consolidation (EWC) (Kirkpatrick et al., 2017) is a technique that regularizes the loss function, such that the model is encouraged to only modify the weights that are less important for the previous tasks. This is achieved by penalizing changes to weights that have large importance for the previously learned tasks, thereby mitigating catastrophic forgetting when learning new tasks.
Memory aware synapses (MAS) (Aljundi et al., 2018) is a regularization-based method that measures the importance of parameters based on the sensitivity of predictions to these parameters. When learning a new task, the model adjusts the weights according to the network’s activations to update the important parameters relevant to the new task data.
Topology-aware weight preserving(TWP) (Liu et al., 2021) is a method introduced for graph continual learning that incorporates weight preservation mechanisms based on graph topology. By considering the local structure of the graph, TWP aims to overcome catastrophic forgetting. After computing the loss, the importance scores of the model weights are calculated according to the topology of the given graph, and the loss is regularized accordingly. This method leverages the properties of the graph.
ER-GNN (Zhou and Cao, 2021) stores samples from previous tasks as experiences and replays them when learning new tasks. After learning task $T_{i}$, sample nodes are saved to the buffer using a selection function. When learning the next task $T_{i+1}$, separate graphs are constructed for each learned task $\{T_{t}\}_{t=1,\dots,i}$. Then, the GNN is trained using these graphs. The overall loss is calculated by regularizing the current task’s loss with the loss from the separately constructed graphs through experience replay.
MetaCLGraph (Unal et al., 2023) combines experience replay and meta-learning. In this method, the initial parameters of the model are calculated using the current task data. When learning a new task, stored samples from previous tasks are merged with the current task data to form a new graph, which is then used to update the model parameters.
FTF-ER (Pang et al., 2024) combines feature and global topological information by normalizing and weightedly integrating two types of node importance scores to evaluate node importance. Specifically, it introduces the Hodge Potential Score (HPS) module to capture global topological information. When learning a new task, a subgraph is induced using all the experience nodes stored in the buffer, the overall loss is calculated as the sum of the loss for the current task and the loss of the subgraph.
B.3. Experimental Setting
@@
B.4. Implementation Detail
To evaluate the effectiveness and generalizability of our method, we conduct experiments on three HGNNs backbones: HAN (Wang et al., 2019), MAGNN (Fu et al., 2020), and HGT (Hu et al., 2020). Adam optimizer is used and the initial learning rate is set to 0.005 for all datasets. For HAN, each task is trained for 200 epochs with an early stopping patience of 100. For MAGNN, each task is trained for 100 epochs, with the early stopping patience set to 5 on DBLP and 10 on IMDB. For HGT, all tasks are trained for 300 epochs with a patience of 30. For datasets trained with mini-batches, the batch size is set to 8. We adopt coverage maximization as the selection function for all experience-replay-based continual learning methods. The buffer size is uniformly set to 50 for the target node type in experience replay baselines (e.g., ER-GNN, MetaCLGraph, and FTF-ER), and 20 under mini-batch settings. For non-target node types, the buffer size is set to 200. The hyperparameters for regularization-based baselines (e.g., EWC, MAS, and TWP) are set following the TWP official repository and the benchmark study (Zhang et al., 2022a).
@@ Appendix C Additional Results and Analysis
C.1. Inference Time Comparison
We provide detailed inference time results in Table 5. We observe that HERO achieves higher efficiency than most baseline methods during the inference phase, further supporting its practicality in real-world deployment.
Table 5. Inference time (s) comparison on four datasets. HERO show the absolute inference time (s), while other methods are reported as relative difference (%). Positive difference means slower than HERO.
Method DBLP IMDB Freebase Yelp
TWP +1.99% +3.81% +2.22% -1.54%
FTF-ER +1.85% +2.98% +2.48% +0.01%
ER-GNN +1.02% +2.78% -1.13% -2.13%
MetaCLGraph +0.13% +3.54% +4.92% +0.80%
HERO (Ours) 3.4124 0.9815 2.5918 9.5556
C.2. Additional Ablation Study